# Attention-Guided Multi-scale Interaction Network for Face Super-Resolution

### v3

@@ Abstract
Recently, CNN and Transformer hybrid networks demonstrated excellent performance in face super-resolution (FSR) tasks. Since numerous features at different scales in hybrid networks, how to fuse these multi-scale features and promote their complementarity is crucial for enhancing FSR. However, existing hybrid network-based FSR methods ignore this, only simply combining the Transformer and CNN. To address this issue, we propose an attention-guided Multi-scale interaction network (AMINet), which contains local and global feature interactions and encoder-decoder phase feature interactions. Specifically, we propose a Local and Global Feature Interaction Module (LGFI) to promote fusions of global features and different receptive fields’ local features extracted by our Residual Depth Feature Extraction Module (RDFE). Additionally, we propose a Selective Kernel Attention Fusion Module (SKAF) to adaptively select fusions of different features within LGFI and encoder-decoder phases. Our above design allows the free flow of multi-scale features from within modules and between the encoder and decoder, which can promote the complementarity of different scale features to enhance FSR. Comprehensive experiments confirm that our method consistently performs well with less computational consumption and faster inference.
@@
Face super-resolution, Hybrid networks, Multi-scale interaction, Attention-guided.I Introduction
@@
Recently, since the advantages exhibited by hybrid networks [5] of CNNs and Transformers in FSR, this type of method has gained increased attention. Specifically, CNN-based FSR methods [6] generally do not require large computational consumption. Still, they specialize in extracting local details, such as the local texture of the face, color, etc., and are unable to model long-range feature interaction, such as the global profile of the face. Transformer-based FSR methods [7] can simulate global modeling well, but their computational consumption is huge. Hybrid FSR methods leverage the strengths of both architectures, enabling efficient extraction of both local and global features while maintaining a reasonable computational overhead. The impressive performance of hybrid-based FSR methods comes from numerous features extracted inside their networks at different scales, such as global features from self-attention, local features from convolution, and features from different stages of the encoder-decoder, which facilitates models to refine local facial details and global facial contours.
@@
However, while existing hybrid-based FSR methods consider utilizing features from different scales to improve FSR, they ignore the problem of how to fuse these multi-scale features to make their properties better complement each other. For example, Faceformer [8] simply parallelizes the connected CNN modules and the window-based Transformer [9] modules. SCTANet [10] also only juxtaposes spatial attention-based residual blocks and multi-head self-attention in designed modules. CTCNet [5] simply connects the CNN module in tandem with the Transformer module. These methods overlook the importance of blending multi-scale features in a complementary manner and enabling smooth information flow across different scales to refine facial details effectively.
@@
To address this problem, we propose an Attention-Guided Multi-scale Interaction Network (AMINet) for FSR in this work. Our AMINet fuses multi-scale features in two main ways, including the fusion of features obtained from self-attention and convolution, and the fusion of features at different stages of the encoder-decoder. Specifically, we design a Local and Global Feature Interaction Module (LGFI) to adaptively fuse global facial and local features with different receptive fields obtained by convolutions. In LGFI, self-attention is responsible for extracting global features, our proposed Residual Depth Feature Extraction Module (RDFE) extracts local features at different scales using separable convolutional kernels of different sizes, and our proposed Selective Kernel Attention Fusion Module (SKAF) is responsible for weighted fusion of these two parts of features for our model to adaptively perform selective fusion during training. In addition, we also utilize our proposed SKAF as a crucial fusion module in our Encoder and Decoder Feature Fusion Module (EDFF) to further perform feature communications of our method by fusing features at different scales from the encoder-decoder processes.
@@
We design an LGFI to differ from the traditional Transformer by allowing free flow and adaptive selective fusion of local and global features within the module.
@@
We design the SKAF to help selective fusions of different scale features within LGFI and EDFF by selecting appropriate convolutional kernels.
In the following sections, we will introduce related work and discuss their disadvantages in Section II. Then, we will introduce our method and proposed core modules in Section III. Next, we will evaluate our method and analyze our ablation experiments in Section IV. Finally, we will discuss our method in Section V and conclude this paper in Section VI.
@@ II-A Face Super-Resolution
Early deep learning approaches focused on leveraging facial priors as guidance to enhance FSR [11] accuracy. For instance, Chen et al. [12] developed an end-to-end prior-based network that utilized facial landmarks and heatmaps to generate FSR images. Similarly, Kim et al. [13] employed a face alignment network for landmark extraction in conjunction with a progressive training technique to produce realistic face images. Ma et al. [14] introduced DICNet, which iteratively integrates facial landmark priors to enhance image quality at each step. Hu et al. [15] explored the use of 3D shape priors to better capture and define sharp facial structures. While these methods have advanced FSR, they require additional labeling of training datasets. Moreover, inaccuracies in prior estimation can significantly diminish FSR performance, especially when dealing with highly blurred face images.
Attention-based FSR methods have been proposed to promote FSR to avoid the adverse effects of inaccurate prior estimates on FSR. Zhang et al. [16] proposed a supervised pixel-by-pixel generation of the adversarial network to improve face recognition performance during FSR. Chen et al. [17] proposed the SPARNet, which can focus on important facial structure features adaptively by using spatial attention in residual blocks. Lu et al. [18] proposed a partial attention mechanism to enhance the consistency of the fidelity of facial detail and facial structure. Bao et al. [19] introduced the equalization texture enhancement module to enhance the facial texture detail through histogram equalization. Wang et al. [20] critical introduced Fourier transform into FSR, fully exploring the correlation between spatial domain features and frequency domain features. Shi et al. [21] designed a two-branch network, which introduces a convolution based on local changes to enhance the ability of the convolution. Liu et al. [10] improves the interaction ability of regional and global features through designed hybrid attention modules. Li et al. [22] designed a wavelet-based network to reduce the loss of downsampling in the encoder-decoder. Although the above methods can reconstruct reasonable FSR images, they cannot promote the efficient fusion of local features with global features and different features at different stages of the encoder-decoder, affecting FSR’s efficiency and accuracy.
@@ II-B Attention-based Super-Resolution
The attention mechanism can improve the super-resolution accuracy of models due to its flexibility in focusing on key areas of facial features. In the super-resolution [23] task, different variants of the attention mechanism include self-attention, spatial attention, channel attention, and hybrid attention.
Zhang et al. [24] inserted channel attention into residual blocks to enhance model representation. Xin et al. [25] utilized channel attention plus residual mechanisms to combine a multi-level information fusion strategy. Chen et al. [17] enhanced FSR by utilizing an improved facial spatial attention that cooperated with the hourglass structure. Gao et al. [26] performed shuffling to hybrid attention. Wang et al. [27] constructed a simplified feed-forward network using spatial attention to reduce parameters and computational complexity. To model long-range feature interaction, the self-attention in Transformer [28, 29] has been widely used in super-resolution. Gao et al. [30] reduced costs by utilizing the recursive mechanism on self-attention. Li et al. [31, 32] combined self-attention and convolutions to complement each other’s required features. Zeng et al. [33] introduced a self-attention network that investigates the relationships among features at various levels. Shi et al. [21] enhanced FSR by mitigating the adverse effects of inaccurate prior estimates through a parallel self-attention mechanism, effectively capturing both local and non-local dependencies. To combine the advantages of different attentions, Yang et al. [34] integrated channel attention with spatial attention to enhance feature acquisition and correlation modeling. Bao et al. [10] and Gao et al. [5] employed spatial attention and self-attention to capture facial structure and details. Zhang et al. [35] employed a hybrid attention that combines self-attention, spatial attention, and channel attention to optimize fine-grained facial details and broad facial structure. Unlike the above attention-based methods that enhance model representation, we utilize attention to learn features from different receptive fields, allowing our network to adaptively select the appropriate convolutional kernel size to match the multi-scale feature fusion. This design enables our network to perform multi-scale feature extraction and improve the integration of features across various scales, leading to enhanced performance and greater adaptability.
@@ III Proposed Method
In this section, we first describe the overall architecture of the proposed Attention-Guided Multi-scale Interaction Network (AMINet). Then, we introduce our local and global feature interaction module (LGFI) and encoder and decoder feature fusion module (EDFF) in our AMINet in detail. Finally, since we provide a GAN version of our AMINet, called AMIGAN, we further introduce the loss functions used to supervise AMIGAN during training.
@@ III-A Overview of AMINet
As illustrated in Fig. 2, our proposed AMINet features a U-shaped CNN-Transformer hierarchical architecture with three distinct stages: encoding, bottleneck, and decoding. For an LR input face image , in the encoding stage, our network aims to extract features at different scales and capture multi-scale feature representations of the input image to get the facial feature . Then, the bottleneck stage network continues to refine the feature and provides a more informative representation to get the refined feature for the subsequent reconstruction phase. In decoding, the network focuses on feature upsample and facial detail reconstruction. Meanwhile, an interactive connection is used between the encoding and decoding stages to ensure the features are fully integrated throughout the network. We can get the reconstructed face feature with rich facial details through the above operators. Finally, through a convolution with reduced channel dimensions plus a residual connection, we get the HR output face image .
@@ III-A3 Decoding stage
In the decoding stage, there are three decoders. We focus on multi-scale feature fusion, aiming at reconstructing high-quality face images at this stage. As depicted in Fig. 2, each decoder includes an upsampling operation, an EDFF, and an LGFI. Each upsampling operator halves the input feature channel counts while doubling the width and weight of the input facial feature. Compared to encoding stages, decoding stages additionally use our proposed SKAF to adaptively and selectively fuse different scale features from the encoder and decoder stages. Through this design, different scale features can interact to recover more detailed face features. The features obtained after three decoders are as follows: , . Finally, a convolution unit is utilized to transform our obtained deep facial feature into an output FSR image .
@@ III-B Local and Global Feature Interaction Module (LGFI)
In our AMINet, LGFI is mainly used for local and global facial feature extraction. As shown in Fig. 2, LGFI consists of Self-attention(SA), a Residual Depth Feature Extraction Module (RDFE), and a Selective Kernel Attention Fusion Module (SKAF), used for local and global feature fusion and interaction, respectively. The SA is designed to extract global features. At the same time, RDEM is designed to extract local features at different scales and enrich local facial details through multiple convolutional kernels under numerous receptive fields. Specifically, the integration of local and global features is achieved through SKAF, which adaptively weights and fuses multi-scale information. SKAF first extracts local and global features via convolutions with different receptive fields and then computes their importance using average pooling. These computed weights are applied to the corresponding features, facilitating an adaptive fusion process. By dynamically selecting and emphasizing key information, SKAF enhances the complementarity between local textures and global structural cues, ultimately improving facial detail reconstruction and overall super-resolution performance.
@@ III-B1 Self-attention (SA)
We utilize Self-attention (SA) to extract global facial features, which can effectively model the relationships between distant features. Meanwhile, through the multi-head mechanism in SR, features can be captured from different subspaces, improving the robustness and generalization ability of the model. As illustrated in Fig. 3 (b), we start by applying a convolutional layer followed by a depth-wise convolutional layer to combine pixel-level cross-channel information and extract channel-level spatial context. From this spatial context, we then generate . For an input facial feature , the process of obtaining can be described as
@@
Next, we reshape , , and into , , and , respectively. After that, the dot product is multiplied by to obtain weights , which facilitates the capturing of the important local context in SA. Finally, we rearrange into . The above operations can be expressed as
@@ III-B2 Residual depth feature extraction module (RDFE)
As shown in Fig. 3 (a), we design RDFE to extract local facial features at different scales. Compared with the traditional feed-forward network (FFN), our RDFE is beneficial for processing more complex features and multi-scale features flexibly. Specifically, for the input feature , we use depthwise convolutions of , , and to parallelly extract three scales facial features, which depthwise convolution can reduce the computational complexity of the model, while convolution with different kernel sizes can effectively extracting rich face details. The reason we employ depthwise separable convolution is to significantly reduce computational complexity while preserving local spatial features. In our model, depthwise convolutions at different scales enable the extraction of fine-grained facial details, such as textures and edge information, across varying receptive fields, enhancing the model’s ability to capture local features efficiently. The above operations can be expressed as
where , , and are , , and depthwise convlution, respectively. However, relying solely on convolutions may lead to insufficient feature fusion across scales, potentially introducing redundant information or affecting the representation of key facial details. To address this issue, we use an attention unit to calculate the feature weight of the fusion feature of three branches. Next, we use the calculated weights to modulate the three-branch features at different scales through element-wise multiplications. This mechanism enables the model to dynamically balance the contributions of multi-scale features, effectively capturing key facial details across varying receptive fields, thereby enhancing detail restoration and overall generalization. To further refine and reconstruct the weighted fused facial features, we apply depthwise convolutions with kernel sizes of , , and , ensuring the joint optimization of local texture details and global structural consistency. This collaborative refinement ultimately leads to higher-quality facial reconstructions. The above operations can be expressed as
where is a concat operator and indicates attention unit. Then, we aggregate the three branches’ features to combine facial detail information under different receptive fields. This process can be described as
where represents convolution. Finally, we utilize a Feature Refinement Module to refine features obtained from previous multiple branches. Specifically, we begin by applying normalization and multiple convolutional layers to refine the local facial context. Afterward, the hourglass block further integrates multi-scale information to capture global and local relationships. The above operations can be expressed as

### v4

# Attention-Guided Multiscale Interaction Network for Face Super-Resolution

@@ Abstract
Recently, CNN and Transformer hybrid networks demonstrated excellent performance in face super-resolution (FSR) tasks. Since numerous features at different scales in hybrid networks, how to fuse these multiscale features and promote their complementarity is crucial for enhancing FSR. However, existing hybrid network-based FSR methods ignore this, only simply combining the Transformer and CNN. To address this issue, we propose an attention-guided Multiscale interaction network (AMINet), which incorporates local and global feature interactions, as well as encoder-decoder phase feature interactions. Specifically, we propose a Local and Global Feature Interaction Module (LGFI) to promote the fusion of global features and the local features extracted from different receptive fields by our Residual Depth Feature Extraction Module (RDFE). Additionally, we propose a Selective Kernel Attention Fusion Module (SKAF) to adaptively select fusions of different features within the LGFI and encoder-decoder phases. Our above design allows the free flow of multiscale features from within modules and between the encoder and decoder, which can promote the complementarity of different scale features to enhance FSR. Comprehensive experiments confirm that our method consistently performs well with less computational consumption and faster inference.
@@
Face super-resolution, Hybrid networks, Multiscale interaction, Attention-guided.I Introduction
@@
Recently, since the advantages exhibited by hybrid networks [5] of CNNs and Transformers in FSR, this type of method has gained increased attention. Specifically, CNN-based FSR methods [6] generally do not require large computational costs. Still, they specialize in extracting local details, such as the local texture of the face, color, etc., and are unable to model long-range feature interaction, such as the global profile of the face. Transformer-based FSR methods [7] can simulate global modeling well, but their computational consumption is huge. Hybrid FSR methods leverage the strengths of both architectures, enabling efficient extraction of both local and global features while maintaining a reasonable computational overhead. The impressive performance of hybrid-based FSR methods comes from numerous features extracted inside their networks at different scales, such as global features from self-attention, local features from convolution, and features from different stages of the encoder-decoder, which facilitates models to refine local details and global contours.
@@
However, while existing hybrid-based FSR methods consider utilizing features from different scales to improve FSR, they ignore the problem of how to fuse these multiscale features to make their properties better complement each other. For example, Faceformer [8] simply parallelizes the connected CNN modules and the window-based Transformer [9] modules. SCTANet [10] also only juxtaposes spatial attention-based residual blocks and multi-head self-attention in designed modules. CTCNet [5] simply connects the CNN module in tandem with the Transformer module. These methods overlook the importance of blending multiscale features in a complementary manner and enabling smooth information flow across different scales to refine facial details effectively.
@@
To address this problem, we propose an Attention-Guided Multiscale Interaction Network (AMINet) for FSR in this work. Our AMINet fuses multiscale features in two main ways, including the fusion of features obtained from self-attention and convolution, and the fusion of features at different stages of the encoder-decoder. Specifically, we design a Local and Global Feature Interaction Module (LGFI) to adaptively fuse global facial and local features with different receptive fields obtained by convolutions. In LGFI, self-attention is responsible for extracting global features, while our Residual Depth Feature Extraction Module (RDFE) extracts local features at different scales using separable convolutional kernels of different sizes, and our Selective Kernel Attention Fusion Module (SKAF) is responsible for weighted fusion of these two parts of features for our model to adaptively perform selective fusion during training. In addition, we also utilize our SKAF as a crucial fusion module in our Encoder and Decoder Feature Fusion Module (EDFF) to further perform feature communications of our method by fusing features at different scales from the encoder-decoder processes.
@@
We design an LGFI to differ from the traditional Transformer by allowing free flow and selective fusion of local and global features within the module.
@@
We design the SKAF to help selective fusions of different-scale features within LGFI and EDFF by selecting appropriate convolutional kernels.
@@ II-A Face Super-Resolution
Early deep learning approaches focused on leveraging facial priors as guidance to enhance FSR accuracy [11]. For instance, Chen et al. [12] developed an end-to-end prior-based network that utilized facial landmarks and heatmaps to generate FSR images. Similarly, Kim et al. [13] employed a face alignment network for landmark extraction in conjunction with a progressive training technique to produce realistic face images. Ma et al. [14] introduced DICNet, which iteratively integrates facial landmark priors to enhance image quality at each step. Hu et al. [15] explored the use of 3D shape priors to better capture and define sharp facial structures. While these methods have advanced FSR, they require additional labeling of training datasets. Moreover, inaccuracies in prior estimation can significantly diminish FSR performance, especially when dealing with highly blurred face images.
Attention-based FSR methods have been proposed to promote FSR to avoid the adverse effects of inaccurate prior estimates on FSR. Zhang et al. [16] proposed a supervised pixel-by-pixel generation of the adversarial network to improve face recognition performance during FSR. Chen et al. [17] proposed the SPARNet, which can focus on important facial structure features adaptively by using spatial attention in residual blocks. Lu et al. [18] proposed a partial attention mechanism to enhance the consistency of the fidelity of facial detail and facial structure. Bao et al. [19] introduced the equalization texture enhancement module to enhance the facial texture detail through histogram equalization. Wang et al. [20] critically introduced the Fourier transform into FSR, fully exploring the correlation between spatial domain features and frequency domain features. Shi et al. [21] designed a two-branch network, which introduces a convolution based on local changes to enhance the ability of the convolution. Liu et al. [10] improves the interaction ability of regional and global features through designed hybrid attention modules. Li et al. [22] designed a wavelet-based network to reduce the loss of downsampling in the encoder-decoder. Although the above methods can reconstruct reasonable FSR images, they cannot promote the efficient fusion of local features with global features and different features at different stages of the encoder-decoder, affecting FSR’s efficiency and accuracy.
@@ II-B Attention-based Super-Resolution
The attention mechanism can improve the super-resolution accuracy of models due to its flexibility in focusing on key areas of facial features. In the super-resolution task, different variants of the attention mechanism include self-attention, spatial attention, channel attention, and hybrid attention [23].
Zhang et al. [24] inserted channel attention into residual blocks to enhance model representation. Xin et al. [25] utilized channel attention plus residual mechanisms to combine a multi-level information fusion strategy. Chen et al. [17] enhanced FSR by utilizing an improved facial spatial attention that cooperated with the hourglass structure. Gao et al. [26] performed shuffling to hybrid attention. Wang et al. [27] constructed a simplified feed-forward network using spatial attention to reduce parameters and computational complexity. To model long-range feature interaction, the self-attention in Transformer [28] has been widely used in super-resolution. Gao et al. [29] reduced costs by utilizing the recursive mechanism on self-attention. Li et al. [30, 31] combined self-attention and convolutions to complement each other’s required features. Zeng et al. [32] introduced a self-attention network that investigates the relationships among features at various levels. Shi et al. [21] enhanced FSR by mitigating the adverse effects of inaccurate prior estimates through a parallel self-attention mechanism, effectively capturing both local and non-local dependencies. To combine the advantages of different attentions, Yang et al. [33] integrated channel attention with spatial attention to enhance feature acquisition and correlation modeling. Bao et al. [10] and Gao et al. [5] employed spatial attention and self-attention to capture facial structure and details. Zhang et al. [34] employed a hybrid attention that combines self-attention, spatial attention, and channel attention to optimize fine-grained facial details and broad facial structure. Unlike the above attention-based methods that enhance model representation, we utilize attention to learn features from different receptive fields, allowing our network to adaptively select the appropriate convolutional kernel size to match the multiscale feature fusion. This design enables our network to perform multiscale feature extraction and improve the integration of features across various scales, leading to enhanced performance and greater adaptability.
@@ III-A Overview of AMINet
As illustrated in Fig. 2, our proposed AMINet features a U-shaped CNN-Transformer hierarchical architecture with three distinct stages: encoding, bottleneck, and decoding. For an LR input face image , in the encoding stage, our network aims to extract features at different scales and capture multiscale feature representations of the input image to get the facial feature . Then, the bottleneck stage network continues to refine the feature and provides a more informative representation to get the refined feature for the subsequent reconstruction phase. In decoding, the network focuses on feature upsample and facial detail reconstruction. Meanwhile, an interactive connection is used between the encoding and decoding stages to ensure the features are fully integrated throughout the network. We can get the reconstructed face feature with rich facial details through the above operators. Finally, through a convolution with reduced channel dimensions plus a residual connection, we get the HR output face image .
@@ III-A3 Decoding stage
In the decoding stage, there are three decoders. We focus on multiscale feature fusion, aiming at reconstructing high-quality face images at this stage. As depicted in Fig. 2, each decoder includes an upsampling operation, an EDFF, and an LGFI. Each upsampling operator halves the input feature channel counts while doubling the width and weight of the input facial feature. Compared to encoding stages, decoding stages additionally use our proposed SKAF to adaptively and selectively fuse different scale features from the encoder and decoder stages. Through this design, different scale features can interact to recover more detailed face features. The features obtained after three decoders are as follows: , . Finally, a convolution unit is utilized to transform our obtained deep facial feature into an output FSR image .
@@ III-B Local and Global Feature Interaction Module (LGFI)
In our AMINet, LGFI is mainly used for local and global facial feature extraction. As shown in Fig. 2, LGFI consists of Self-attention(SA), a Residual Depth Feature Extraction Module (RDFE), and a Selective Kernel Attention Fusion Module (SKAF), used for local and global feature fusion and interaction, respectively. The SA is designed to extract global features. At the same time, RDEM is designed to extract local features at different scales and enrich local facial details through multiple convolutional kernels under numerous receptive fields. Specifically, the integration of local and global features is achieved through SKAF, which adaptively weights and fuses multiscale information. SKAF first extracts local and global features via convolutions with different receptive fields and then computes their importance using average pooling. These computed weights are applied to the corresponding features, facilitating an adaptive fusion process. By dynamically selecting and emphasizing key information, SKAF enhances the complementarity between local textures and global structural cues, ultimately improving facial detail reconstruction and overall super-resolution performance.
@@ III-B1 Self-attention (SA)
We utilize Self-attention (SA) to extract global facial features, which can effectively model the relationships between distant features. Meanwhile, through the multi-head mechanism in SR, features can be captured from different subspaces, improving the robustness and generalization ability of the model. As illustrated in Fig. 3 (b), we start by applying a convolutional layer followed by a depth-wise convolutional layer to combine pixel-level cross-channel information and extract channel-level spatial context. From this spatial context, we then generate . For an input facial feature , the process of obtaining can be described as:
@@
Next, we reshape , , and into , , and , respectively. After that, the dot product is multiplied by to obtain weights , which facilitates the capturing of the important local context in SA. Finally, we rearrange into . The above operations can be expressed as:
@@ III-B2 Residual depth feature extraction module (RDFE)
As shown in Fig. 3 (a), we design RDFE to extract local facial features at different scales. Compared with the traditional feed-forward network (FFN), our RDFE is beneficial for processing more complex features and multiscale features flexibly. Specifically, for the input feature , we use depthwise convolutions of , , and to parallelly extract three scales facial features, which depthwise convolution can reduce the computational complexity of the model, while convolution with different kernel sizes can effectively extracting rich face details. The reason we employ depthwise separable convolution is to significantly reduce computational complexity while preserving local spatial features. In our model, depthwise convolutions at different scales enable the extraction of fine-grained facial details, such as textures and edge information, across varying receptive fields, enhancing the model’s ability to capture local features efficiently. The above operations can be expressed as:
where , , and are , , and depthwise convlution, respectively. However, relying solely on convolutions may lead to insufficient feature fusion across scales, potentially introducing redundant information or affecting the representation of key facial details. To address this issue, we use an attention unit to calculate the feature weight of the fusion feature of three branches. Next, we use the calculated weights to modulate the three-branch features at different scales through element-wise multiplications. This mechanism enables the model to dynamically balance the contributions of multiscale features, effectively capturing key facial details across varying receptive fields, thereby enhancing detail restoration and overall generalization. To further refine and reconstruct the weighted fused facial features, we apply depthwise convolutions with kernel sizes of , , and , ensuring the joint optimization of local texture details and global structural consistency. This collaborative refinement ultimately leads to higher-quality facial reconstructions. The above operations can be expressed as:
where is a concat operator and indicates attention unit. Then, we aggregate the three branches’ features to combine facial detail information under different receptive fields. This process can be described as:
where represents convolution. Finally, we utilize a Feature Refinement Module to refine features obtained from previous multiple branches. Specifically, we begin by applying normalization and multiple convolutional layers to refine the local facial context. Afterward, the hourglass block further integrates multiscale information to capture global and local relationships: